Company A has overstated the fair value of net assets of its acquisition of Company B in 2017. The price of acquisition was 169. Company A estimated goodwill to be 53. In 2018, you as an analyst want to make an accounting adjustment to the recorded fair value net assets by writing-down 100% of them based on newly acquired information. However, the goodwill will remain on the balance sheet. With what amount will the shareholder’s equity in 2018 decrease (in absolute terms) based on this adjustment? Consider that the tax rate is 20%.
Answer #1

Goodwill = Price of acquisition - Fair value of Net assets

Thus, Fair value of net assets = 169 - 53 = 116

Value of net assets written down by 116 i.e. 100%

Tax benefit of such write down = 116 * 20% = 23.20

Thus, net decrease in shareholder's equity = (116 -23.20) = 92.80